# High-throughput electro-optic upconversion and downconversion with few-photon added noise

### v1

@@ Abstract
A transducer bridging microwave and optical frequencies with sufficiently low noise would allow the deterministic generation of arbitrary optical quantum states for applications in metrology, the storage and manipulation of optical quantum information, and the networking of remote superconducting quantum processors. Reducing input-referred added noise $N_{\text{add}}$ to as low a level as possible is imperative for such quantum state manipulation. At the same time, it is useful to increase the rate of transduction to enable transducing signals at rates that are compatible with the lifetimes of logical qubits that process the transduced quantum information. Current state-of-the-art lifetimes in superconducting architectures are approximately 1 ms using physical qubits [1, 2], 1 ms using error-corrected bosonic encodings [3], and 300 $\mu$s using logical encodings of many physical qubits [4]. Finite qubit lifetimes place constraints not only on the transducer bandwidth $B$ (which must be fast enough to transduce quantum information within the time qubits can faithfully store it), but also on the overall signal throughput in order to support an entangled bit rate much faster than decoherence rates. Signal throughput is further reduced by less-than-unity transducer efficiency $\eta$ and duty cycle $D$, underscoring the importance of increasing the throughput of quantum transducers, even as progress is made on logical lifetimes of quantum memories. In practice, there is often a tradeoff between noise performance and throughput, as increasing throughput frequently involves increasing the average drive power, which can lead to heating-induced noise.
@@
Concurrent with improving the performance of electro-optic transducers, a fully developed transducer platform would be capable of operating in both directions of frequency conversion. For instance, there are several proposed protocols for sharing entanglement between separated quantum systems, and these place different demands on the transducers in a network of microwave-frequency nodes linked optically. Long-range quantum networks dominated by loss will likely benefit from using heralding protocols, in which entanglement infidelity due to inefficiency is mitigated at the cost of transmission rates and greater noise sensitivity [5]. Such a protocol favors a network topology with transducers operating in upconversion (from microwave to optical frequencies) to take advantage of mature optical single-photon-detection technologies [6]. On the other hand, transducers operating in downconversion (from optical to microwave frequencies) would be useful for generating entanglement from squeezed optical light [7]. Further, deterministic state transfer protocols [8, 9] are possible if efficiencies are high, and require one transducer operating in upconversion and one in downconversion. In general, a transducer’s performance can be asymmetric between the two directions of frequency conversion, depending on the origin of the dominant source of noise. Published work to date has typically operated more favorably in upconversion than in downconversion, in the relatively few works where downconversion performance has been demonstrated and quantified [10, 11].
@@
In this work, we present measurements of a transducer with high throughput, $\Theta=\eta BD$ exceeding 7 kHz, and noise performance close to input-referred added noise $N_{\text{add}}=1$, expressed in units of photons, or photons per second per Hz of bandwidth, in upconversion ($N_{\text{add}}^{\text{up}}=2.6$) as well as downconversion ($N_{\text{add}}^{\text{down}}=3$). While similar throughputs have been achieved in a device with quantum-enabled performance in upconversion [12], in downconversion, our device increases throughput by three orders of magnitude over the leading quantum demonstration [11], at the cost of a single order of magnitude worse noise performance. Our platform achieves such performance due to its robustness to strong optical pumping, as demonstrated in previous work [13], as well as significant improvement to the electrical circuit that we show here.
@@
$N_{\text{add}}^{\text{up}}(\omega_{\text{m}})$ $=\underbrace{\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{e}}}+n_{\text{em}}+\frac{n_{\text{om}}\Gamma_{\text{o}}}{\Gamma_{\text{e}}}}_{N_{\text{add,m}}}+\underbrace{\frac{\bar{n}_{\text{o}}\Gamma_{\text{T}}^{2}}{\Gamma_{\text{e}}\Gamma_{\text{o}}}}_{N_{\text{add,e\kern-0.50003pt/\kern-0.74997pto}}}-\underbrace{\frac{2\bar{n}_{\text{o}}\Gamma_{\text{T}}}{\Gamma_{\text{e}}}}_{N_{\text{add,corr}}}$ (1)
@@
$N_{\text{add}}^{\text{down}}(\omega_{\text{m}})$ $=\underbrace{\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{o}}}+n_{\text{om}}+\frac{n_{\text{em}}\Gamma_{\text{e}}}{\Gamma_{\text{o}}}}_{N_{\text{add,m}}}+\underbrace{\frac{\bar{n}_{\text{e}}\Gamma_{\text{T}}^{2}}{\Gamma_{\text{o}}\Gamma_{\text{e}}}}_{N_{\text{add,e\kern-0.50003pt/\kern-0.74997pto}}}-\underbrace{\frac{2\bar{n}_{\text{e}}\Gamma_{\text{T}}}{\Gamma_{\text{o}}}}_{N_{\text{add,corr}}},$ (2)
where $n_{\text{th}}$ is the environmental mechanical occupation, $\bar{n}_{\text{o}}$ is the effective occupation of the optical cavity due to thermal occupation or the presence of technical noise on the optical pump, and $n_{\text{om}}=\bar{n}_{\text{o}}+n_{\text{min,o}}$ is the bath occupation that the mechanical mode is coupled to optomechanically, with $n_{\text{min,o}}$ denoting the optomechanical backaction limit [24]. The quantities associated with the electrical port $n_{\text{em}}$, $\bar{n}_{\text{e}}$, and $n_{\text{min,e}}$ are defined analogously. In the resolved-sideband limit, $n_{\text{om}}=\bar{n}_{\text{o}}$, but we differentiate these noise contributions explicitly here to clarify their physical origins. In Eqs. 1 and 2, the first three terms comprise noise from motion of the transduction mode $N_{\text{add,m}}$, and the fourth term $N_{\text{add,e\kern-0.70004pt/\kern-1.04996pto}}$ is from noise purely of electromagnetic origin measured at the output port. The fifth term $N_{\text{add,corr}}$ results from destructive interference between the fields responsible for the third and fourth terms.
@@
If motion of the membrane were the only source of noise, like it is in upconversion, our device would demonstrate quite asymmetrical performance, favoring operation in downconversion with large $\Gamma_{\text{o}}$. However, when our device operates in downconversion, the microwave field also contributes to $N_{\text{add}}$ through $N_{\text{add,e\kern-0.70004pt/\kern-1.04996pto}}$. Assuming $\Gamma_{\text{T}}=\Gamma_{\text{o}}+\Gamma_{\text{e}}$ to combine the last three terms of Eq. 2, we have
@@
In Figs. 3(c) and (d), we put the respective performance measurements in upconversion and downconversion in the context of other results in the field, using a format similar to that first used in Ref. [12]. We see that the throughputs we achieve at the noise levels we measure are consistent with the best achieved in the field in upconversion, although unlike the device in Ref. [12], the turnaround due to thermomechanical motion at low throughputs prevents our current device from achieving quantum operation. Relative to our previous work [13], this turnaround occurs at higher throughput because of a $\gamma_{\text{m}}$ that is over an order-of-magnitude higher, as measured electromechanically by ringdown of the transduction mode. In downconversion, our platform achieves over three orders of magnitude higher throughput, at the cost of one order of magnitude increase in $N_{\text{add}}$, compared with the leading downconversion demonstration in the field [11].
@@
Reasonable improvements to our device design would allow quantum-enabled noise performance. Refinement of the engineered mechanical isolation of our pinned membrane would suppress the stiff modes in the mechanical spectrum and address the elevated mechanical intrinsic loss rate measured in this device (Methods). Recovering the mechanical coherence we have measured in previous generations of this device, combined with parallel improvement to the electromechanical design to increase the electromechanical coupling rate [36], would yield high-throughput quantum-enabled performance in both upconversion and downconversion. With transduction demonstrations improving in both throughput and added noise, we are rapidly approaching transducer integrated quantum capacities at the kHz-scale, at which point transducers would support, on average, the entanglement of one pair of superconducting qubits within every 1 ms physical lifetime.

### v2

@@ Abstract
[…]
@@
Concurrent with improving the performance of electro-optic transducers, a fully developed transducer platform would be capable of operating in both directions of frequency conversion. For instance, there are several proposed protocols for sharing entanglement between separated quantum systems, and these place different demands on the transducers in a network of microwave-frequency nodes linked optically. Long-range quantum networks dominated by loss will likely benefit from using heralding protocols, in which entanglement infidelity due to inefficiency is mitigated at the cost of transmission rates and greater noise sensitivity [5]. Such a protocol favors a network topology with transducers operating in upconversion (from microwave to optical frequencies) to take advantage of mature optical single-photon-detection technologies [6]. On the other hand, transducers operating in downconversion (from optical to microwave frequencies) would be useful for generating entanglement from squeezed optical light [7]. Further, deterministic state transfer protocols [8, 9] are possible if efficiencies are high, and require one transducer operating in upconversion and one in downconversion. In general, a transducer’s performance can be asymmetric between the two directions of frequency conversion, depending on the origin of the dominant source of noise. Published work to date has typically operated more favorably in upconversion than in downconversion, and there are few works in which downconversion performance has been quantified [10].
@@
In this work, we present measurements of a transducer with high throughput, $\Theta=\eta BD$ exceeding 7 kHz, and noise performance close to input-referred added noise $N_{\text{add}}=1$, expressed in units of photons, or photons per second per Hz of bandwidth, in upconversion ($N_{\text{add}}^{\text{up}}=2.6$) as well as downconversion ($N_{\text{add}}^{\text{down}}=3$). While similar throughputs have been achieved in a device with quantum-enabled performance in upconversion [11], in downconversion, our device increases throughput by nearly four orders of magnitude over previous work at the threshold of quantum-enabled performance [10]. Our platform achieves such performance due to its robustness to strong optical pumping, as demonstrated in previous work [12], as well as significant improvement to the electrical circuit that we show here.
@@
$N_{\text{add}}^{\text{up}}(\omega_{\text{m}})$ $=\underbrace{\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{e}}}+n_{\text{em}}+\frac{n_{\text{om}}\Gamma_{\text{o}}}{\Gamma_{\text{e}}}}_{N_{\text{add,m}}}+\underbrace{\frac{\bar{n}_{\text{o}}\Gamma_{\text{T}}^{2}}{\Gamma_{\text{e}}\Gamma_{\text{o}}}}_{N_{\text{add,e\kern-0.6806pt/\kern-1.0208pto}}}-\underbrace{\frac{2\bar{n}_{\text{o}}\Gamma_{\text{T}}}{\Gamma_{\text{e}}}}_{N_{\text{add,corr}}}$ (1)
@@
$N_{\text{add}}^{\text{down}}(\omega_{\text{m}})$ $=\underbrace{\frac{n_{\text{th}}\gamma_{\text{m}}}{\Gamma_{\text{o}}}+n_{\text{om}}+\frac{n_{\text{em}}\Gamma_{\text{e}}}{\Gamma_{\text{o}}}}_{N_{\text{add,m}}}+\underbrace{\frac{\bar{n}_{\text{e}}\Gamma_{\text{T}}^{2}}{\Gamma_{\text{o}}\Gamma_{\text{e}}}}_{N_{\text{add,e\kern-0.6806pt/\kern-1.0208pto}}}-\underbrace{\frac{2\bar{n}_{\text{e}}\Gamma_{\text{T}}}{\Gamma_{\text{o}}}}_{N_{\text{add,corr}}},$ (2)
where $n_{\text{th}}$ is the environmental mechanical occupation, $\bar{n}_{\text{o}}$ is the effective occupation of the optical cavity due to thermal occupation or the presence of technical noise on the optical pump, and $n_{\text{om}}=\bar{n}_{\text{o}}+n_{\text{min,o}}$ is the bath occupation that the mechanical mode is coupled to optomechanically, with $n_{\text{min,o}}$ denoting the optomechanical backaction limit [23]. The quantities associated with the electrical port $n_{\text{em}}$, $\bar{n}_{\text{e}}$, and $n_{\text{min,e}}$ are defined analogously. In the resolved-sideband limit, $n_{\text{om}}=\bar{n}_{\text{o}}$, but we differentiate these noise contributions explicitly here to clarify their physical origins. In Eqs. 1 and 2, the first three terms comprise noise from motion of the transduction mode $N_{\text{add,m}}$, and the fourth term $N_{\text{add,e\kern-0.79727pt/\kern-1.19579pto}}$ is from noise purely of electromagnetic origin measured at the output port. The fifth term $N_{\text{add,corr}}$ results from destructive interference between the fields responsible for the third and fourth terms.
@@
If motion of the membrane were the only source of noise, like it is in upconversion, our device would demonstrate quite asymmetrical performance, favoring operation in downconversion with large $\Gamma_{\text{o}}$. However, when our device operates in downconversion, the microwave field also contributes to $N_{\text{add}}$ through $N_{\text{add,e\kern-0.79727pt/\kern-1.19579pto}}$. Assuming $\Gamma_{\text{T}}=\Gamma_{\text{o}}+\Gamma_{\text{e}}$ to combine the last three terms of Eq. 2, we have
@@
In Figs. 3(c) and (d), we put the respective performance measurements in upconversion and downconversion in the context of other results in the field, using a format similar to that first used in Ref. [11]. We see that the throughputs we achieve at the noise levels we measure are consistent with the best achieved in the field in upconversion, although unlike the device in Ref. [11], the turnaround due to thermomechanical motion at low throughputs prevents our current device from achieving quantum operation. Relative to our previous work [12], this turnaround occurs at higher throughput because of a $\gamma_{\text{m}}$ that is over an order-of-magnitude higher, as measured electromechanically by ringdown of the transduction mode.
In downconversion, our device demonstrates a leap in achieved performance, with an increase in throughput of more than a factor of 5400 compared with previous work [10]. This orders-of-magnitude improvement is accompanied by only a factor-of-three greater $N_{\text{add}}$.
@@
Reasonable improvements to our device design would allow quantum-enabled noise performance. Refinement of the engineered mechanical isolation of our pinned membrane would suppress the stiff modes in the mechanical spectrum and address the elevated mechanical intrinsic loss rate measured in this device (Methods). Recovering the mechanical coherence we have measured in previous generations of this device, combined with parallel improvement to the electromechanical design to increase the electromechanical coupling rate [36], would yield high-throughput quantum-enabled performance in both upconversion and downconversion. With transduction demonstrations improving in both throughput and added noise, we are approaching transducer integrated quantum capacities at the few-kHz scale, at which point transducers would support, on average, the entanglement of a pair of superconducting qubits within every 1 ms physical lifetime.